A florist brought 240 rose at rs 9 per dozen .if he sold all of them at rs 1 each ,what profit did he make ? Also what is his profit per cent ?

Grade:9

2 Answers

aneesh
8 years ago
60 is profit and 33 1/3% is his profit
 
RISHABH CHAURASIA
8 years ago
C.P of 1 rose = 9/12 = 0.75S.P of 1 rose = Re 1Hence, since S.P. is greater than C.P. so he has made profit.Profit = S.P. – C.P. = 1-0.75 = 0.25Hence, profit on sale of 1 rose = 0.25This means that the profit on sale of 240 roses = 240 x 0.25 = Rs 60Profit % = (Profit/C.P.) x 100= (0.25/0.75) x 100= 100/3